OpenStudy (anonymous):

A student has mean score of 88 on five test taken. What score must she obtain on her next test to have a mean (average) score of 80 on all six test?

OpenStudy (anonymous):

88*5= 440 (440+ x)/6= 80 x= 40

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

soory how did you get 40

OpenStudy (anonymous):

440+x = 80*6 440 + x= 480 x= 40 Average = total scores / 6 tests the totals would be 440 + the unknown score, which I named x

OpenStudy (anonymous):

Got it thanks!